OpenStudy (anonymous):

Through (−8, −6); perpendicular to the line passing through (−5, 6) and (−1, 4)

OpenStudy (ranga):

Find the slope of the line passing through (−5, 6) and (−1, 4) m = (y2-y1)/(x2-x1) ---- (1) Slope of perpendicular line is: - 1 / m So y = -1/m(x) + b ----- (2) (put m from (1)) The perpendicular line passes through (-8, -6). Put x = -8 and y = -6 above in (2) and solve for b.

OpenStudy (anonymous):

thank can I please get a little more details

OpenStudy (ranga):

Slope of the line m passing (−5, 6) and (−1, 4) = (y2-y1)/(x2-x1) m = (4-6)/(-1-(-5)) = -2/(-1+5) = -2/4 = -1/2 Slope of the perpendicular line = -1/m = -1/(-1/2) = 2 y = mx + b = 2x + b when x = -8, y = -6 -6 = 2(-8) + b -6 = -16 + b b = 16-6 = 10 y = 2x + 10

OpenStudy (anonymous):

thank you

OpenStudy (ranga):

you are welcome.